a 0.582-mol sample of an ideal gas is at 15*c and 81.8 kpa. calculate the volume for the ideal gas, as described by the ideal gas low.
a.0.887 L
b.10.4 L
c.17.0 L
d.22.4 L
SHOW ALL WORK
1
Expert's answer
2017-04-20T06:57:17-0400

Answer:
pV = nRT
R = 8.314 L · kPa · K-1 · mol-1
81.8 · V = 0.582 · 8.314 · (273+15)
V = (0.582 · 8.314 · (273+15))/81.8 = 17.0 L
Answer c) 17.0 L.